### v1

Abstract
Extended exposure to virtual reality environments can induce motion sickness, often referred to as cybersickness, which may lead to physiological stress responses and impaired cognitive performance. This study investigates the aftereffects of VR-induced motion sickness with a focus on physiological stress markers and working memory performance. Using a carousel simulation to elicit cybersickness, we assessed subjective discomfort (SSQ, FMS), physiological stress (salivary cortisol, alpha-amylase, electrodermal activity, heart rate), and cognitive performance (n-Back task) over a 90-minute post-exposure period. Our findings demonstrate a significant increase in both subjective and physiological stress indicators following VR exposure, accompanied by a decline in working memory performance. Notably, delayed symptom progression was observed in a substantial proportion of participants, with some reporting peak symptoms up to 90 minutes post-stimulation. Salivary cortisol levels remained elevated throughout the observation period, indicating prolonged stress recovery. These results highlight the need for longer washout phases in XR research and raise safety concerns for professional applications involving post-exposure task performance.
keywords:
Virtual Reality, Cybersickness, Stress Response, Working Memory, Human Factors1 Introduction
Virtual Reality (VR) and Augmented Reality (AR) technologies are increasingly being integrated into cognitively demanding work environments, such as aviation, surgery, and complex manufacturing processes. These Extended Reality (XR) technologies offer immersive training opportunities, real-time data visualization, and enhanced decision-making support. However, despite its advantages, prolonged exposure to XR can introduce physiological and cognitive challenges, including motion sickness, fatigue, and stress responses, which may impact performance and well-being [52, 57]. Understanding how these environments affect users is crucial for optimizing their application in high-stakes professional settings.
A key issue in immersive environments is cybersickness, a form of motion sickness induced by sensory discrepancies between visual and vestibular signals [53]. It manifests as nausea, dizziness, headache, and other autonomic symptoms frequently experienced by VR users [10, 31]. Motion sickness in VR can occur under various conditions, including prolonged exposure, high visual flow, latency between head movements and visual updates, and discrepancies between acceleration cues and expected motion [68]. The severity and duration of symptoms vary among individuals, with some experiencing mild discomfort that subsides within minutes, while others report persistent symptoms lasting hours after VR exposure [39]. Factors such as individual susceptibility, prior VR experience, and display technology play significant roles in determining the intensity of motion sickness effects.
The washout phase of sickness after XR exposure has been studied, though not in great depth. However, it remains a critical aspect to explore further, particularly given its implications for both the robustness of scientific methods [85] and safety in high-stakes environments. For instance, understanding how long symptoms persist and how quickly they subside is essential when considering the use of XR technologies in professions where precise performance is required, such as surgery or aviation. Prolonged exposure to immersive technologies may lead to lingering symptoms that could impair a user’s ability to operate potentially dangerous machinery or perform high-risk tasks. As XR technologies become more prevalent in work environments, studying recovery from sickness is vital to ensure safe and effective usage in critical settings.
Previous studies have demonstrated that cybersickness can elicit a stress response, reflected in autonomic and endocrine changes. Research has shown that individuals experiencing higher discomfort in VR environments exhibit increased cortisol levels, altered Electrodermal Activity (EDA), and Heart Rate (HR) patterns [6, 50]. A recent study by [34] further supports this by demonstrating that cybersickness caused by Head-Mounted Display (HMD)-based VR environments significantly influences physiological responses, including elevated Heart Rate Variability (HRV), altered autonomic nervous system activity, and increased cortisol levels. The physiological stress response is characterized by the activation of the Hypothalamic-Pituitary-Adrenal (HPA) axis and the sympathetic nervous system [46]. While the HPA axis regulates cortisol release, the sympathetic nervous system governs rapid autonomic adaptations, manifesting in increased EDA and changes in HR [76].
Stress has been shown to impact cognitive functions, particularly working memory. While acute stress can enhance cognitive performance under specific conditions [58], extensive research suggests that high stress levels impair working memory performance [60, 64]. This impairment may be due to increased cortisol and norepinephrine secretion, which modulate neural networks in the prefrontal cortex, thereby affecting working memory function [4].
Based on these findings, the present study examines the effects of motion sickness in a VR environment on physiological and cognitive parameters. Specifically, we investigate how VR-induced motion sickness progresses after stimulation and whether it elicits a physiological stress response, and whether this response affects working memory performance. We hypothesize that VR stimulation (carousel ride) leads to a heightened physiological stress response compared to the resting condition, as measured by cortisol and alpha-amylase concentrations, EDA, and HR. Furthermore, we predict that cognitive performance in the working memory task (n-Back task) is significantly reduced following VR stimulation compared to the resting condition.
This study extends previous research by incorporating multiple physiological markers, allowing for a more comprehensive understanding of the motion sickness stress response in VR environments. By examining these effects in a controlled experimental setting, this research contributes to a deeper understanding of the interaction between VR, motion sickness, stress, and cognitive performance.
2 Related Work
2.1 Temporal Progression of Cybersickness During Exposure
Existing research offers diverse perspectives on the temporal dynamics of cybersickness symptoms during exposure to virtual environments.
Several studies indicate that cybersickness tends to increase progressively with prolonged exposure. [29] observed a linear rise in symptoms among fighter pilots exposed to VR conditions over three hours. Similarly, [63] reported an upward trend in symptoms over a 16-minute duration, while [56] found a comparable linear increase within the first ten minutes. [47] noted a similar trajectory in symptoms over approximately six minutes. [84], who conducted self-reported assessments across two 15-minute conditions, also found indications of a continuous increase.
Longer-duration studies reinforce this trend. [69] examined exposure durations of 15, 30, 45, and 60 minutes in a between-subjects design, identifying a sustained increase in cybersickness beyond the 15-minute mark. [38] further corroborated this finding, showing that longer exposures correlated with more intense symptoms.
A meta-analysis by [59] categorized studies into exposure groups of 0-10 minutes, 10-20 minutes, and beyond 20 minutes. Their findings indicated stable oculomotor symptoms across these groups, whereas nausea and disorientation peaked between 10 and 20 minutes. However, this contrasts with other research suggesting a sustained increase in symptoms beyond 20 minutes.
Some studies suggest adaptation effects over prolonged exposure. In a two-hour experiment, symptoms plateaued after 75 minutes, hinting at potential habituation [9]. [48] found no significant Simulator Sickness Questionnaire [28] (SSQ) score differences between early and later exposure stages, reinforcing the notion of early adaptation. Similarly, studies using the Fast Motion Sickness Questionnaire [30] (FMS) values [47, 74] primarily within the first ten minutes reported adaptation trends.
[67] investigated cybersickness in treadmill-based virtual experiences, noting a significant increase in SSQ disorientation scores up to 15 minutes, followed by an adaptation phase at 45 minutes. Interestingly, nausea peaked at 45 minutes, aligning with [21]’s findings, whereas oculomotor symptoms plateaued after 15 minutes. Their data suggest a more complex progression where adaptation occurs alongside symptom escalation.
Other research challenges the notion of a purely gradual increase in symptoms, instead proposing an episodic pattern [79, 81, 73]. Such fluctuations may be difficult to detect due to limitations in sampling rates, which are often constrained by practical feasibility.
Taken together, these findings highlight both a general trend of increasing cybersickness over time and potential adaptation effects at various stages of exposure, emphasizing the need for finer-grained temporal analyses to capture the nuanced progression of symptoms.
2.2 Temporal Progression of Cybersickness Post Stimulation
Research is sparse on how cybersickness reacts post-stimulation and shows significant interindividual differences in this aspect. For instance, while most subjects experience minimal symptoms after a 10-60 minute break, some continue to suffer from severe symptoms [66].
A key issue in cybersickness research is the lack of uniformity in post-exposure studies. The need for standard methodologies in this area has been discussed, with suggestions for establishing specific standards [14].
In a study measuring SSQ 30 minutes after VR exposure, no significant differences were found compared to a pre-study baseline, suggesting rapid recovery in most subjects. However, the disorientation subscale remained significantly higher than pre-experiment levels [66]. Contrastingly, [17]’s study indicates more prolonged effects of cybersickness. They measured cybersickness before, immediately after, and one, two, and three hours post-exposure over three days. SSQ values were significantly higher post-exposure each day, challenging the notion of rapid recovery suggested by Singer et al. Yet, both studies found significantly lower SSQ values in the first follow-up compared to immediately post-exposure. Further, recovery patterns vary with the VR experiment’s format. In experiments with high SSQ scores, some participants needed over 30 minutes to recover. In contrast, others reported no discomfort just five minutes post-experiment [72]. This observation was mirrored by [71], where most of the participants no longer felt any symptoms 40 minutes after exposure, whereas 14% still felt strong after-effects.
Few studies have examined participants’ well-being the day after the experiment. For example, a study with pilots post-simulator training revealed that 4.6% reported symptoms 24 hours after training [77]. Another study with pilots showed that those with intense symptoms had significantly higher values two hours post-experiment compared to a less affected group, but no differences were observed after six, 12-18, and 24 hours [70].
In summary, cybersickness symptoms typically last 1-2 hours but can persist for days in extreme cases [24, 54, 75, 81]. Additionally, numerous studies consistently demonstrate an association between habituation or adaptation to cybersickness and repeated exposure to VR applications [24, 54, 75, 81]. There is much less knowledge about the precise models governing the decay over time.
2.3 Cybersickness and Biophysical Responses
Traditionally, subjective measures such as the SSQ and the FMS have been used to assess the severity of cybersickness. However, these self-reported tools have limitations, including relying on users’ subjective perception, which can be influenced by factors such as mood, prior VR experience, and individual sensitivity. Consequently, there has been growing interest in the use of biophysical measures as objective indicators of cybersickness.
Biophysical measures, such as Electroencephalography (EEG), Electrocardiography (ECG), EDA, and HRV, offer the potential to complement or even replace subjective measures. These physiological signals provide direct insights into the autonomic nervous system’s response to VR stimuli, making them valuable for more accurately assessing the onset and severity of cybersickness. Additionally, biophysical signals may enable the development of real-time systems that predict and mitigate cybersickness, offering a more personalized and precise approach to user comfort.
Recent studies have demonstrated that these biophysical measures can effectively capture the physiological responses associated with cybersickness. For instance, changes in HR, skin conductivity, and stomach activity have been shown to correlate with the severity of cybersickness symptoms [10, 16]. Similarly, EEG measures, particularly changes in brainwave patterns, have been identified as reliable indicators of cybersickness [40, 41, 33]. These findings suggest that biophysical measures could effectively supplement subjective measures like the SSQ, offering more accurate, continuous, and objective assessments of cyber sickness.
[55] employed a combination of EEG, Electromyography (EMG) from the gastrocnemius muscles, and HR to assess motion sickness, finding that EMG parameters were more significant than EEG for predicting motion sickness. In a similar vein, [42] demonstrated that higher power in the gamma band (32Hz) of EEG, coupled with increased heart rate, was associated with heightened cybersickness severity.
The role of multiple physiological signals in predicting cybersickness has been explored in several studies. [10] utilized Electromyography (EOG), EDA, Electrooculography (EGG), ECG, and respiration to study cybersickness. They found that changes in stomach activity, blinking behavior, and breathing patterns could predict cybersickness severity. Likewise, [16] collected data on ECG, EOG, respiratory signals, and skin conductivity from 66 participants and showed that these physiological signals were valuable indicators of cybersickness.
[25] compared HR and EDA signals between high and low cybersickness groups, finding significant differences in HR but no significant differences in breathing patterns or HRV. These results suggest that HR is a more reliable indicator of cybersickness than other measures like respiration rate or HRV.
[40] proposed a method for recognizing VR motion sickness using EEG rhythm energy ratios, achieving high recognition accuracy for individual subjects using a wavelet packet transform. Their method, which combines EEG features with machine learning models such as support vector machines, highlights the growing trend of using computational models to improve the accuracy of motion sickness detection.
[45] explored how HRV and EDA metrics during VR tasks could be used as potential indicators of cybersickness. Similarly, [11] employed a multi-modal approach, incorporating EEG, ECG, EGG, and postural sway to predict motion sickness, finding that EEG-derived features were the most effective in cybersickness prediction, with an accuracy of 93.8%.
More recently, [26] introduced a multimodal deep fusion network that combined HR, EDA, eye movements, and head-tracking data to predict the onset of cybersickness in real-time. These advances reflect a broader trend in the field where most research focuses on ensemble methods and deep learning models to predict cybersickness onset, rather than evaluating and understanding individual contributions of biophysical signals. Similarly, [51] proposed machine learning models using physiological and behavioral features, such as heart rate and body movements, achieving promising results in predicting cybersickness onset.
[34] simulated aircraft motion in VR, and physiological responses were measured, including heart rate, blood pressure, body temperature, and cortisol levels. Their results showed significant increases in heart rate and cortisol levels associated with cybersickness, while blood pressure decreased. The study concludes that HR and cortisol are promising physiological markers for cybersickness assessment and calls for future research on factors specifically affecting disorientation.
In conclusion, while there is substantial research on the relationship between biophysical responses and cybersickness, most studies concentrate on using machine learning techniques, particularly ensemble models, to predict the onset of cybersickness. The integration of multiple biosignals offers promising results, but the individual contributions of each signal remain underexplored, and further research is needed to investigate specifically post-exposure physiological responses.
3 Method
To address our research questions, particularly regarding the progression of cybersickness symptoms after stimulation, we deliberately induced cybersickness in participants (study condition) and subsequently monitored them using physiological and endocrinological measures, established subjective scales such as the SSQ and FMS, and a cognitive load task. Participants were observed for up to 90 minutes post-stimulation.
To normalize individual differences and properly contextualize, we paired this design with a sequence-balanced within-subject control condition in which no sickness stimulus was applied. The study was reviewed and approved by the university’s ethics board. It was conducted in line with the Declaration of Helsinki.
The following sections provide a detailed overview of the study methodology. The study design was preregistered via OSF.
3.1 Stimulus (IV) & Apparatus
To induce a cybersickness stimulus, we utilized an open-access simulator developed by [82]. This simulator immerses users in the perspective of a seated passenger in a cart on an amusement ride.
It is designed to provide a more controlled stimulus than typical virtual roller coasters (cf. [3]), while still maintaining a realistic and fun experience, as opposed to highly abstract approaches such as an optoelectronic drum (cf. [8]).
Throughout the experience, participants remained physically seated. There was no interaction beyond the ability to look around freely. Their only task was to experience and observe the ride passively. The simulator is modeled after a real amusement ride known as the Brakedancer and is implemented in Unity. The base platform of the simulator is tilted by 8 degrees. We used no virtual body representation in the study. Each cart is grouped with three others around a central pivot point. Individual carts can rotate around their axes, specifically the yaw and pitch axes. Additionally, all cart groups on the tilted platform collectively rotate around the platform’s central axis, which is aligned with a central tower structure.
At the start of the simulation, all three rotational axes begin moving at a slow pace. To ensure participant safety, the experiment was equipped with a dead man’s switch, allowing the simulation to be halted at any time if a participant reported reaching 70% of their cybersickness threshold (see Section 3.3.1).
The amusement ride is set within a virtual fairground environment, which includes typical elements such as a free-fall tower, various booths, and a Ferris wheel. During the simulation, both the participant’s cart and the surrounding gondolas begin to rotate, gradually increasing in speed according to a predefined automatic temporal protocol (every 30 seconds). The complete parameterization is accessible with the publicly available source code. The experience is presented from a first-person perspective in VR. To deliver the stimulus, we employed an HTC VIVE Pro headset.
3.2 Procedure
Participants were informed during recruitment that the experiment aimed to induce cybersickness. Each participant was invited to two afternoon sessions scheduled at the same time but several days apart. This timing was crucial for endocrinological measurements, as hormone levels exhibit significant fluctuations throughout the day. For the same reason, participants were instructed to refrain from physical exercise, alcohol, caffeine, and medications at least 24 hours before testing, as well as to refrain from consuming anything but water two hours prior to testing. Additionally, participants agreed to hand in their mobile phones for the duration of the two-hour experiment to prevent uncontrolled influences on their stress levels, such as those caused by accessing news or messages.
Upon arrival at the first session, participants received both written and verbal explanations of the experiment. They then signed an informed consent form before being randomly assigned to either the control condition, which did not include a cybersickness stimulus, or the study condition, in which the stimulus was presented (see Section 3.1). A schematic representation of the experimental procedure is shown in Fig 2.
3.2.1 Study Condition
On the study condition day of the experiment, the participant’s interpupillary distance was measured using the GlassesOn app and adjusted in the HMD to ensure an accurate VR presentation. Subsequently, participants completed the SSQ for the first time, assessing their current symptoms related to cybersickness. The experimenter then introduced the FMS scale, explaining that a score of 0 represented no symptoms, while a score of 10 indicated severe sickness, such as nausea leading to vomiting. Participants were informed that during the sickness stimulus, they would be asked every 30 seconds to report their current sickness level via FMS. If they reported a score of 7 or higher, the experiment would be stopped immediately, and they were made aware of this. We did not aim to maximize sickness in this study, as participants needed to remain able to continue with the experiment and subsequent observations after the stimulus.
After introducing participants to the stimulus, the SSQ, and the FMS, the experimenter collected the first saliva sample (, where denotes the start of the stimulation). Following this, physiological measurements were prepared. Participants were fitted with a chest strap to monitor respiration, and ECG electrodes were placed in a triangular formation on both clavicles and the lower left side of the thoracic cage. The electrode sites were cleaned with alcohol (propanol) before placement. Electrodes for measuring EDA were attached to the participant’s non-dominant hand and secured with tape. These physiological measurements were only recorded in the experimental condition and served as control variables.
To conclude the preparation phase, the first baseline psychophysiological activity was recorded for three minutes, followed by a second saliva baseline measurement (). Participants were then introduced to the VR environment, and after familiarization, another psychophysiological baseline was recorded for 3 minutes. Before starting the stimulation at , the experimenter reiterated the termination criterion: a self-reported FMS score of 7 or higher.
The VR stimulation lasted either for a maximum of eight minutes or until the participant reached the termination threshold. After the stimulus ended either way, participants remained in the VR environment for an additional minute and were then asked to report their FMS score. The HMD was then removed, followed by the final 3-minute psychophysiological measurement—the recovery phase—and another saliva sample ().
After completing this phase, participants moved to a separate resting room. There, additional saliva samples, SSQ, and FMS assessments were collected at predefined time points (, , , and ). At , participants completed an n-Back task. The resting room was designed to be a low-stimulation environment, providing a quiet space where participants could read pre-selected magazines or their own literature.
3.2.2 Control Condition
The control condition is particularly important for the endocrinological markers, as these are susceptible to environmental influences. Moreover, it became evident that subjective measurements such as the SSQ do not fully adhere to the zero-baseline assumption [7]. As a result, these measurements also become more robust when an appropriate control condition is implemented. We opted for a low-level (passive) control condition rather than closely mimicking the experimental condition. This decision was made to control for the well-known circadian rhythm of the HPA axis [35]. Participants in the control group were immediately directed to the resting room, where saliva samples and SSQ assessments were taken at the same time points as in the experimental condition. The first sample was labeled , following the notation of the study condition. They also completed the n-Back task at .
3.3 Measures (DV)
3.3.1 Questionnaires
Simulator Sickness Questionnaire (SSQ)
The participants filled out the Simulator Sickness Questionnaire [28] (SSQ) at six time points, once before the experiment () and five times following the experiment (, , , , and ). All timestamps refer to the cortisol measurement, while the SSQ assessment always took place immediately before it, with a time offset of approximately 2 minutes. This instrument measures the symptoms related to cybersickness of nausea, disorientation, and oculomotor problems on a 4-point Likert scale from None to Severe using 16 items. It provides the three subscales of nausea, disorientation, and oculomotor as well as an overall score of cybersickness.
Fast Motion Sickness Scale (FMS)
Further, the participants were asked to rate their feelings of comfort throughout the study condition on a scale from 0 (not sick at all) to 10 (frank sickness). This scale is an adapted version of the Fast Motion Sickness Questionnaire [30] (FMS). It was adapted from 0 to 20 to make it more intuitive for the participants under the potentially stressful circumstances this experiment evokes (cf. [83, 1]). The FMS was used during the experiment to evaluate the current well-being of the participants during the stimulation. For this purpose, participants were asked to rate their current sickness every 30 seconds during the motion stimulation. Once they reached the threshold of 7, the stimulation was terminated to prevent the participants from severe physical reactions such as throwing up. One minute after the experiment, as well as with every SSQ following (study condition only), the participants gave another rating on the FMS.
Visually Induced Motion Sickness Susceptibility Questionnaire (VIMSSQ)
Furthermore, the participants were asked to fill out the Visually Induced Motion Sickness Susceptibility Questionnaire [32] (VIMSSQ) in the control condition between and . By using this instrument, we expect to evaluate the susceptibility of the participants to show symptoms related to motion sickness while using visual displays. It is expected to predict the SSQ values that the participants provide following the cybersickness-inducing experiment.
Demographic Questionnaire
We prepared a demographic questionnaire, asking the participants about their sex, age, weight, height, physical activities, their experience in 3D video games, their experience in VR environments, their subjective susceptibility to motion sickness, how they respond to stress as well as whether they expect to experience cybersickness.
3.3.2 Salivary Cortisol and Alpha-Amylase
Saliva samples were collected at seven time points throughout the experiment (, , , , , , ). The samples were collected by the participants themselves under supervision and with instructions on the proper procedure. They were given a cotton swab, which they moved around the inside of their mouth for 45–60 seconds, focusing particularly on the cheeks. Afterward, they placed the swab into the designated and labeled tube and handed it over to the experimenter. The samples were initially stored at the university facility where the experiment was conducted and later transported to the laboratory. At both locations, they were preserved at -18°C until biochemical analysis.
For this experiment, we used Salivettes (Sarstedt, Nümbrecht, Germany). Salivary cortisol levels were measured using a time-resolved fluorescence immunoassay [13]. Additionally, alpha-amylase activity was determined using 2-chloro-4-nitrophenyl-D-maltotriose (CNP-G3) as the relevant substrate.
The enzymatic action of alpha-amylase on this substrate yields 2-chloro-p-nitrophenol, which can be spectrophotometrically measured at 405 nm. The amount of alpha-amylase activity present in the sample is directly proportional to the increase in absorbance at 405 nm [44, 80]. The intra-assay Coefficient of Variation (CV) was 4.0-6.7%, the corresponding inter-assay CV was 7.1-9.0% for cortisol 2.8 - 6.3% (intra), and 5.5 – 7.6% (inter) for sAA. Analyses were performed in duplicate and averaged.
3.3.3 Psychophysiological Measures
All psychophysiological measures were recorded between approximately and minutes of the study condition with an integrated multichannel amplifier (Brainproducts V-Amp 16, Gilching, Germany) attached to a standard PC. Signals and section markers were recorded with Brainvision Recorder software (Brainproducts, Gilching, Germany) at a sampling rate of 1000 Hz and exported to EDF files for further preprocessing and analyses.
Heart Rate (HR) & Heart Rate Variability (HRV)
To measure Heart Rate (HR) and calculate the Heart Rate Variability (HRV) throughout the experiment, a single-channel ECG was recorded. Disposable electrodes were attached beneath the right collarbone and below the left ribcage (cf. Einthoven II). Peak detection and semiautomatic artifact rejection were done with Kubios HRV Scientific (Kubios, Kuopio, Finland). Kubios was also used for the segmentation and calculation of average HR and HRV RMSSD for the baseline, motion, and recovery periods.
Electrodermal Activity (EDA)
The Electrodermal Activity (EDA), more specific the Skin Conductance Level, was recorded from the index and ring finger of the non-dominant hand using Ag/AgCl electrodes (4 mm diameter) filled with a conductive paste for EDA recordings (Neurospec AG, Stans, Switzerland). This raw signal was low-pass filtered (0.3 Hz, Butterworth, 4th order), and segmented and averaged (median) to calculate SCL for the baseline, motion, and recovery period.
3.3.4 n-Back Task
Participants have completed a series of n-Back tasks at in both conditions. The n-Back task consisted of three blocks with ascending working memory load (1-back to 3-back). In each block, 60 items (14 different capital letters, 15 target stimuli) were presented. Each trial started with a blank screen (1000 ms) followed by a letter (max. 2000 ms until response or time out). For each stimulus presented, participants had the task of indicating by keystroke whether it had been presented one, two, or three times previously or not. The stimuli were presented centrally on a PC screen (specify), and the responses were recorded using a standard PC keyboard with two color-coded keys. Each block lasted about 3 minutes. The presentation was made using PsychoPy (insert current version and citation). Correct vs. incorrect responses and reaction times were recorded for subsequent statistical analysis.
3.4 Participants
According to [34], we expected a medium effect size of in the psycho-endocrinological responses. To calculate the required sample size, we used G*power 3.1 [15], which resulted in a minimum of participants (, ). By listing this experiment on our intern study list as well as advertising through the E-Mail newsletter of the university, we were able to recruit the necessary amount of participants. Participants who smoked excessively ( cigarettes a day) (cf. [36, 5]), were pregnant, or took cortisol-containing medication were excluded from the study. Thirty-five healthy participants were enrolled in the study.
Some participants’ datasets had to be removed during data analysis due to incompleteness and other issues. For example, one participant was excluded because their saliva sample during the control condition was a statistically significant outlier. This participant appeared to be exhausted and stressed, which aligns with the statistical analysis. The final sample comprised participants (male/female: 15/15), aged 19-44 years (; ).
Each participant gave informed consent and was compensated 60€for participating or had the chance to acquire study credits. The ethics committee of Trier University approved this study.
4 Results
Two-factorial repeated-measure ANOVAs were conducted to test for differences in self-reported motion sickness (SSQ), FMS scores, salivary cortisol, and salivary alpha-amylase in both conditions (study vs. control) and over the sampling point (time). For HR, HRV, and EDA, one-factorial ANOVAs were calculated to test for linear and quadratic effects between baseline, motion, and recovery periods. Effect sizes for all ANOVAs are reported as partial . To explain variance regarding the psychophysiological response within the study condition, separate multiple regression analyses were calculated with the predictors: age, sex, termination of motion, and SSQ. All statistical analyses were done in IBM SPSS Statistics for Mac OS (Version 30.0). All figures were created in GraphPad Prism (version 10.4.1, San Diego, CA). The level of significance was set to .
4.1 Subjective Responses
Simulator Sickness Questionnaire (SSQ)
The SSQ scores over the course of the two conditions are depicted in Fig. 3a. The repeated measures ANOVA revealed a significant main effect of condition, indicating that the SSQ significantly differed between the study and control conditions. Specifically, participants in the study condition exhibited significantly higher simulator sickness scores compared to those in the control condition (, , ). A significant main effect of time was also observed, demonstrating that simulator sickness scores varied across the six experimental time points (, , ). Additionally, the interaction between condition and time was statistically significant (, , ), suggesting that the effect of motion on simulator sickness evolved.
Differential condition effects were also shown to a similar extent for the three SSQ subscales nausea (, , ), oculomotor discomfort (, , ) and disorientation (, , ) - see Fig. 3b.
Fast Motion Sickness (FMS)
In addition to the repeated assessments of the SSQ, motion sickness was assessed with the FMS (exclusively in the study condition). Throughout the stimulation, the subjects were asked for a rating every 30 seconds and, on average, showed the expected increase. At an intensity of 7 (on a scale of 0-10), the stimulation was terminated (see Section 3.3.1). Following the stimulation, the values then returned to the initial level after 45 minutes at the latest, as expected (, , ) - see Fig. 3c.
These findings indicate that exposure to motion leads to a progressive increase in simulator sickness, which peaks during the exposure as indicated by the FMS ratings and develops differently compared to the control condition as indicated by the SSQ ratings over the whole experiment (see Fig. 3).
4.2 Salivary Cortisol and Alpha-Amylase
The salivary cortisol and alpha amylase levels over the course of the two conditions are depicted in Fig. 4.
Salivary Cortisol Levels
The repeated measures ANOVA revealed a significant main effect of condition, indicating that motion sickness had a statistically significant impact on salivary cortisol levels. Specifically, the analysis showed that participants exhibited significantly higher cortisol levels in the study condition compared to the control condition (, , ). A significant main effect of the sample was observed, demonstrating that cortisol levels varied across different samples (, , ). Additionally, the interaction between condition and sample was statistically significant (, , ), indicating that the effect of motion sickness on cortisol levels differed depending on the sample. These findings suggest that exposure to motion sickness induces a pronounced HPA axis response, as reflected in elevated cortisol levels (see Fig. 4a). In the study condition, the average response magnitude was (). Seventy percent of the participants showed an increase of at least and could thus be classified as responders (cf. [49, 37]).
Alpha Amylase
The repeated measures ANOVA also revealed a significant main effect of condition for salivary alpha-amylase levels, indicating that motion sickness had a statistically significant impact (, , ). Participants exhibited significantly higher alpha-amylase levels in the study condition compared to the control condition. A significant main effect of the sample was found (, , ), suggesting that alpha-amylase levels varied over time. Furthermore, the interaction between condition and sample was statistically significant (, , ), indicating that the effect of motion sickness on alpha-amylase levels depended on the sample. These results suggest that motion sickness elicits a measurable adrenergic response, as reflected in increased alpha-amylase levels (see Fig. 4b).
4.3 Physiological Measures
| Baseline | Motion | Recovery | ANOVA | Quadratic Contrast | ||||||
| Heart rate (bpm) (n=28) | 80.0 | 10.4 | 86.5 | 12.0 | 79.2 | 11.0 | 30.10 | .001 | 71.79 | .001 |
| HRV (RMSSD) (n=28) | 38.1 | 15.6 | 32.4 | 14.5 | 40.2 | 22.0 | 8.48 | .001 | 24.19 | .001 |
| Skin conductance level (n=26) | 3.98 | 2.82 | 5.00 | 3.47 | 5.01 | 3.70 | 9.70 | .001 | 14.14 | .001 |
Within-subject effects of the sickness stimulus on HR, HRV, and EDA were tested using separate repeated measures ANOVAs with a planned quadratic contrast to specifically test for effects during motion compared to baseline and recovery periods. Descriptives and results of the ANOVAs are found in Fig. 5 and Table 1.
In sum, the stimulation significantly increased HR, increased EDA, and significantly decreased HRV (as indexed by the RMSSD), indicating that the study condition induced measurable activation of the cardiovascular system, autonomic activation, and specifically a reduction in parasympathetic tone as indicated by the effect on HRV RMSSD.
In all three measures, the quadratic contrast suggested u-shaped effects over the experiment, however, skin conductance did not decrease during the recovery period.
4.4 Working Memory Performance
Performance in the n-Back task was tested based on error rates and reaction times. Descriptive values are found in Fig. 6.
Errors in the n-Back task
The ANOVA revealed a significant main effect of condition, indicating that VR-induced motion sickness had a statistically significant impact on the number of errors made. Specifically, the study condition compared to the control condition (, , ). A highly significant main effect of cognitive load was observed, demonstrating that increasing cognitive load strongly influenced task performance (, , ). Participants made more errors as the cognitive load increased, confirming that higher task difficulty impaired performance. The interaction between condition and cognitive load was not statistically significant (, , ). These findings suggest that while both VR-induced motion sickness and cognitive load independently affect cognitive performance, the effects of motion sickness become more pronounced as task difficulty increases.
Reaction times (RT)
For Reaction Time (RT), the ANOVAs results showed a significant main effect of condition, , , , indicating that participants in the study condition exhibited longer RT compared to those in the control condition. Additionally, there was a strong and highly significant main effect of cognitive load, , , , with RTs increasing as cognitive load intensified. However, the interaction between condition and load did not reach statistical significance, , , , suggesting that the effect of cybersickness on RTs remained stable across different levels of cognitive load. These findings indicate that cybersickness leads to prolonged reaction times, and higher cognitive load further increases RTs, but the impact of motion sickness does not vary significantly depending on load level.
4.5 Predictors of the Physiological Response to Motion in VR
| Cortisol () | |||
| Constant | 1.525 | .140 | |
| Sex | -0.295 | -1.694 | .103 |
| Age | -0.022 | -0.126 | .901 |
| Terminated | 0.404 | 2.109 | .045* |
| SSQ | -0.299 | -1.555 | .132 |
| a-Amylase () | |||
| Constant | 0.037 | .971 | |
| Sex | 0.128 | 0.690 | .497 |
| Age | 0.080 | 0.431 | .670 |
| Terminated | 0.204 | 0.996 | .329 |
| SSQ | -0.381 | -1.818 | .081 |
| Heart rate () | |||
| Constant | -0.196 | .846 | |
| Sex | 0.267 | 1.398 | .176 |
| Age | 0.152 | 0.796 | .434 |
| Terminated | -0.280 | -1.351 | .190 |
| SSQ | 0.192 | 0.926 | .364 |
| HRV () | |||
| Constant | -0.907 | .374 | |
| Sex | 0.018 | 0.103 | .919 |
| Age | 0.017 | 0.100 | .921 |
| Terminated | 0.509 | 2.712 | .012* |
| SSQ | -0.503 | -2.677 | .013* |
| Skin conductance () | |||
| Constant | 0.569 | .575 | |
| Sex | 0.142 | 0.715 | .482 |
| Age | 0.061 | 0.304 | .764 |
| Terminated | -0.294 | -1.346 | .192 |
| SSQ | -0.085 | -0.389 | .701 |
To explore potential linear associations between age, sex, termination of the VR exposure, SSQ and physiological responses to sickness stimulation in VR, we conducted separate multiple regression analyses for each physiological measure, using the individual peak-baseline differences of the respective physiological measure as the criterion (see Table 2). Overall, the only significant predictor in all analyses was the variable that coded the termination of motion in VR by the participant for the regression on cortisol. All other predictors were not significant. To illustrate the selective effect, we plotted the cortisol response over the experiment for those who terminated the VR exposure and those who were exposed the full 8 minutes – see Fig. 7. The corresponding mixed ANOVA revealed a significant main effect of time (, , ), a significant main effect of group (, , ), and a significant interaction (, , ), indicating pronounced cortisol responses in the participants who terminated the VR exposure before the regular end after .
5 Discussion
5.1 SSQ Progression and Implications
First, we must acknowledge that the instrument used—the so-called Cybersicker—effectively induced cybersickness as intended. We observed a solid response compared to the baseline in both subjective measures, such as the SSQ and FMS, and objective measures, including cortisol, alpha-amylase, and, ultimately, the n-Back task. Fifteen minutes after the onset of stimulation, the average SSQ score in the study condition reached (), indicating a significantly elevated level (compared to , in the control condition). As expected, the results lie within a moderate range, also when compared to other studies (see [65] for a systematic overview). For direct comparison, [34] are cited again due to the close scientific alignment, reporting a total SSQ score of () pre-stimulation and () post-stimulation; note that no control condition was included in their study. This aligns with our design goal, which aimed to induce a level of cybersickness that still allowed participants to be systematically observed and interviewed after exposure. It is plausible that the immediate post-exposure response was even higher. However, we deliberately chose to standardize the measurement time across all participants relative to the onset of the stimulus rather than its conclusion. Together with the necessary follow-up for biophysical measurements and VR offboarding, this resulted in a delay of the first SSQ measurement post-stimulation by approximately 7 to 12 minutes, as previously mentioned—depending on whether and when the session was terminated. The shortest exposure was recorded for a subject who reached an FMS score of 7 in just under 3 minutes. Therefore, any statements regarding the period immediately following the exposure and the first measurement remain speculative.
When examining the SSQ subscales, the strongest response is found in the Nausea and Disorientation scales, consistent with other studies using exposure durations comparable to the one applied here [59].
5.1.1 Delayed Symptoms
One of the motivations for this study and the post-hoc analysis of cybersickness exposure was the recurring anecdotal reports of delayed symptoms—such as headaches—that appear in some individuals with a considerable time lag after exposure [81]. This phenomenon is particularly relevant when considering safety aspects of VR usage (see Section 5.3), but it also raises important methodological concerns. Delayed symptom onset could compromise the validity of within-subject study designs by introducing uncontrolled, higher-order carry-over effects [85].
Indeed, we observed such delayed reactions in the SSQ scores of several participants. Five out of thirty participants did not reach their peak SSQ score immediately after exposure, but only at a later measurement point (2 at 30 min, 1 at 45 min, 1 at 60 min, and 1 at 90 min). A closer look at the subscales reveals that these delayed effects occur almost exclusively in the Oculomotor and Disorientation subscales—and even more prominently so. In total, 9 participants reached their peak value in one of these scales only after the first post-stimulus measurement (3 at 30 min, 3 at 45 min, 1 at 60 min, and 2 at 90 min for Oculomotor; and 3 at 30 min, 3 at 45 min, and 1 at 90 min for Disorientation). In summary, 10 out of 30 participants showed delayed symptoms, represented in at least one of these subscales.
In contrast, the Nausea subscale showed the strongest initial increase but consistently decreased afterward—monotonically in all but one case—indicating that it was not affected by delayed effects. As a result, the overall SSQ trend may obscure lingering symptoms in the other two subscales. These findings provide empirical support—possibly for the first time—for the existence of significantly delayed symptoms, with some effects peaking up to 90 minutes or later post-exposure. This observation urgently calls for further investigation.
It should also be noted that, due to the study design, we cannot rule out the possibility that the n-Back task administered after the 30-minute SSQ measurement may have interacted with or influenced the reported symptoms at 45, 60, and 90 minutes (cf. [62]). Replication of the results under controlled variation of this factor is therefore necessary.
5.1.2 Time of Recovery
Another notable observation is that, on average, participants returned to the SSQ baseline level of the control condition after 90 minutes. This is a considerable duration, aligning with the few long-term observations reported in related studies (see Section 2.2) and raising concerns about washout periods in within-subjects experiments, which are typically much shorter. However, it is important to note that the n-Back task, as just mentioned, conducted at minutes after stimulation onset (approximately 30 minutes post-stimulation), may have interacted with participants’ SSQ score through items such as fullness of head. On the other hand, the fact that our stimulation protocol resulted in 70% nausea-induced early terminations suggests that recovery times could be even longer under stronger stimulation conditions.
5.2 Cortisol Reaction - Cybersicker as a Stressor?
Based on findings from previous studies, we expected a stress response including self-reported stress, elevated salivary cortisol as well as autonomic activation [34]. Overall, we found the expected pattern of increased subjective stress and physiological responses. However, given that cortisol is primarily considered a marker for social stress [12], the average magnitude of the baseline-to-peak response—approximately 300% of baseline—is surprisingly high and has rarely been found in stress induction studies, neither with an in-person settings [78], nor in VR setups [86, 43]. However, it should be taken into account that the individual responses in the present study vary significantly. Taking this high variability into account, we found a baseline-peak effect size of Cohen’s , which is roughly comparable to previous studies with psychological stressors (such as the TSST) and previous studies inducing cybersickness.
For instance, [34] observed a cortisol level before cybersickness induction of () and after stimulation of (), corresponding to an average increase of 37%. Due to the lower variability, this corresponds to an effect size of Cohen’s . In comparison, meta-analyses summarizing TSST studies reported overall pre-post effect sizes of approximately Hedges - [20, 61], while studies using VR versions of the TSST observed slightly lower effect sizes around [22].
The cortisol response to cybersickness induction opens up possibilities for using the Cybersicker as a stressor in future stress research. This might be especially interesting since cybersickness, stress, and psychological stress might be based on different mechanisms. While psychological stress presumably stimulates the HPA axis by inducing a perceived threat to the psychological self, rotation-associated nausea in VR might rather represent a direct physiological threat to the individual’s physical integrity. However, further investigations are necessary, as our study design does not rule out the possibility that the stress response was amplified by psychological factors, such as the presence of the experimenter triggering participants to feel more embarrassed about experiencing nausea in the presence of others. A direct comparison of the stress response to psychological stress and cybersickness within the same sample in future studies could help to better understand the hypothesized differences or similarities between the two stressors.
Furthermore, we observed a delayed response of salivary cortisol compared to the SSQ, with peaks occurring at and after the start of the simulation. This is expected, as cortisol is released with a temporal delay relative to the stressor and can be measured in saliva with an additional delay [20]. However, 90 minutes after the experiment began, cortisol levels had still not returned to control levels and remained almost three times higher than baseline. This extended recovery period may be due to the strong stress response (with a pronounced cortisol increase) in our study or related to other aspects of the Cybersicker like physiological stress responses in other systems, e.g., the cardiovascular system [34, 2].
5.3 n-Back Task and Implications for Safety with and after VR usage
In the present study, we observed that regardless of the cognitive load of the working memory task, the induction of cybersickness led to a significant drop in performance compared to the control condition. One explanation for the reduced working memory performance after cybersickness induction relates to the well-established effects of delayed effects of stress-induced cortisol secretion on working memory [19, 18]. Presumably, elevated cortisol levels approximately 30 minutes after stress induction might have impaired working memory performance, as reflected in reduced accuracy on the n-Back task compared to the control condition. This finding aligns with previous research demonstrating that acute stress and the subsequent glucocorticoid response can transiently disrupt prefrontal cortex functioning, which is critical for working memory processes [64]. Cortisol’s delayed effects may interfere with neural activity in the dorsolateral prefrontal cortex, leading to diminished cognitive flexibility and attentional control [4].
These results support the notion that the temporal dynamics of the cybersickness-induced stress response are crucial for understanding the temporal transition of potential VR-related cognitive impairments. In addition, our results suggest that even approximately 30 minutes after stimulation, cybersickness continues to have a significant negative impact on cognitive processes. This is further supported by subjective assessments using the SSQ, indicating that participants, upon prompted self-reflection, are aware that they are still experiencing the aftereffects of cybersickness. If these two measures did not align, it would have far-reaching safety implications.
However, our experimental setup does not allow us to determine whether subjective discomfort and cognitive performance return to baseline simultaneously, necessitating further investigations. Nevertheless, we emphasize that cybersickness has a clearly measurable impact on well-defined cognitive variables even after exposure, which, to our knowledge, is the first time this has been explicitly demonstrated. This highlights the importance of seriously considering cybersickness and its potential aftereffects, particularly in safety-critical contexts such as operating vehicles, machinery, or other high-risk tasks in VR environments.
5.4 Onset of Sickness & Correlations
Although the stimulation phase was not the primary focus of this investigation, several physiological signals were recorded for control purposes. These are briefly discussed below.
The FMS scores increased throughout sickness stimulation, but the effect appeared to diminish toward the end of the exposure phase, suggesting a non-linear trajectory (see Fig. 3). Physiological measures followed the general trend of a reaction: HR increased during stimulation, while HRV decreased, both returning to baseline levels during the recovery phase. EDA also increased with stimulation but exhibited a more delayed reaction, remaining elevated throughout the recovery period. This general pattern is consistent with prior findings in real-world settings and VR research [10, 6, 50].
The SSQ response was significantly predicted only by HRV (, , ) and a binary variable coding early termination of the stimulation phase before reaching the maximum exposure duration of 8 minutes (, , ). No other physiological variable showed a significant relationship with acute SSQ scores. The VIMSSQ was only predicting the HR response (), and showed no significant associations with any other physiological variables.
Concerning endocrinological response (salivary cortisol), the only significant predictor in all regression models was the variable coding voluntary termination of VR motion by the participant. No other predictors—including HR, HRV, or EDA—were statistically significant. This outcome is not necessarily unexpected and aligns with earlier findings, especially when considering the relatively small sample size (), which limits the detection of subtle effects such as those associated with sex differences [27, 23].
5.5 Limitations
In addition to the limitations already discussed, our study design presents several constraints that should be considered.
A low-level (passive) control condition was chosen to account for the well-known circadian rhythm of the HPA axis [35]. Consequently, it cannot be ruled out that part of the observed response was influenced by unspecific contextual factors of the test situation, such as the VR environment or the presence of the experimenter. Future studies should incorporate active control conditions to better quantify the specific effects of HPA activation induced by rotational movement in VR.
Furthermore, our stimulation lasted a maximum of eight minutes and was not designed to push participants to their individual limits. This implies that participants may not have experienced the full spectrum of cybersickness symptoms. A meta-analysis by [59] categorized studies into exposure durations of 0–10 minutes, 10–20 minutes, and beyond 20 minutes. Their findings indicated that while oculomotor symptoms remained stable across these groups, nausea and disorientation peaked between 10 and 20 minutes. This suggests that our results may not be fully generalizable to longer exposure durations or more intense sickness conditions.
6 Conclusion
This study provides compelling evidence that VR-induced motion sickness elicits not only immediate discomfort but also sustained physiological stress responses and cognitive impairments. The consistent elevation of salivary cortisol and the delayed peak in subjective symptoms suggest that recovery from cybersickness is both protracted and complex. Our results underscore the importance of considering delayed and lingering effects in both experimental designs and applied VR use cases, particularly in high-stakes professional environments. Furthermore, the observed reduction in working memory performance following sickness induction highlights the potential safety implications when VR is used prior to cognitively demanding tasks. Future research should aim to disentangle the contributions of physiological and psychological stress components in cybersickness and develop guidelines for safe VR deployment in real-world settings.
Competing Interests
The authors declare that they have no financial or non-financial conflicts of interest related to the subject matter or materials discussed in this manuscript. No funding was received for the conduct of this study.
Compliance with Ethical Standards
This study involved human participants and was conducted in accordance with the ethical principles outlined in the Declaration of Helsinki. Informed consent was obtained from all participants prior to their involvement. The study protocol was reviewed and approved by the local ethics committee.